OpenStudy (anonymous):

What is the volume of the box if it is scaled down by a factor of $1/10$? I dont get this question. But the Volume is 132. So what do I do with it?

OpenStudy (anonymous):

@january123

OpenStudy (anonymous):

so anyone?

OpenStudy (january123):

Sorry, im not on the right device...i would have googled the question...try that

OpenStudy (anonymous):

i cant google it though the volume answer is from another question...

OpenStudy (anonymous):

@GretaKnows do u know?

OpenStudy (anonymous):

@SyedaLovesPie

OpenStudy (anonymous):

@pooja195 @TheSmartOne

OpenStudy (anonymous):

@Austin6i6

pooja195 (pooja195):

@TheSmartOne GET OVER HERE >:(

TheSmartOne (thesmartone):

I'm here. OS takes a while to lag :P

OpenStudy (anonymous):

so can u help me @TheSmartOne ?

TheSmartOne (thesmartone):

So what is the formual for volume?

OpenStudy (anonymous):

the formula is V=whl

OpenStudy (anonymous):

in this case V=4*3*11=132

OpenStudy (anonymous):

@TheSmartOne

TheSmartOne (thesmartone):

Ok, so we are reducing it by a scale factor if 1/10. So that means the new volume will be: \(\sf\Large\frac{4}{10}\times\frac{3}{10}\times\frac{11}{10}\)

OpenStudy (anonymous):

oh ok that would equal 0.132

TheSmartOne (thesmartone):

And basically that means that your answer is: \(\sf\large\frac{132}{10^3}\) It's 10^3 because we are dealing with volume. If the area was reduced by 1/10 then it would be 10^2 in the denominator.

TheSmartOne (thesmartone):

And good job! :)

OpenStudy (anonymous):

oh ok thank you!!!!!!!!

TheSmartOne (thesmartone):

No problem. :)

TheSmartOne (thesmartone):

\(\color{blue}{\text{Originally Posted by}}\) @pooja195 @TheSmartOne GET OVER HERE >:( \(\color{blue}{\text{End of Quote}}\) Oh, I did get over here :P